### v1

@@ 2.3 Extracted frequencies
Figure 1: Stacked representation of the extracted independent, significant TESS frequencies for 212 of the Gaia discovered $\beta$ Cep stars, sorted by the TESS frequency with the largest amplitude. Most of the stars exhibit multi-periodic behaviour with excited modes over a large frequency range. The grey-scale gives the amplitudes relative to the largest amplitude of each star on a logarithmic scale. This figure includes only the stars whose main frequency occurs in the plotted $\beta$ Cep range. Some stars have lower main frequencies and turn out to be SPB stars or $\beta$ Cep/SPB hybrid pulsators with dominant high-order g-modes, or are dominated by other variability than pulsations.
To analyse a large number of pulsators, stacking and sorting the periodograms can reveal hidden structures (e.g. Bowman & Kurtz, 2018; Li et al., 2020b; Hey & Aerts, 2024). Figure 1 shows the extracted significant frequencies from the TESS photometry for our Gaia-observed sample sorted by the frequency of highest amplitude. Since $\beta$ Cep stars show typically few excited low-order modes that do not occur in an asymptotic regime of high or low frequencies, no obvious structure emerges from this diagram. Nevertheless, this visual representation of the sample shows the range of detected significant frequencies. We observe that the distribution flattens towards low frequencies while declining more abruptly towards high frequencies. This is fully in line with instability computations of low-order modes in young and evolved $\beta$ Cep stars (Pamyatnykh, 1999). However, the sharp cutoff near 8 d-1 may also be partly due to the natural transition towards the $\delta$ Sct star frequency regime adopted in the classification algorithms (Gaia Collaboration et al., 2023). Further, it becomes obvious from the figure that the observed significant frequencies for a given star depend on the position of its dominant frequency.
Figure 2: Comparison between the Gaia and TESS frequencies for 215 stars. Top: TESS frequency $f_{\mathrm{TESS}}$ against the Gaia derived frequency $f_{\emph{Gaia}}$. Black symbols indicate stars for which we found a match in frequency, red symbols show stars with matching frequencies for other than the dominant signal, and blue symbols indicate stars without a counterpart signal in the TESS data. Lines are connecting the offset frequencies to their matched counter part. For three stars the main frequency is below the plotting range. Bottom: Difference between the Gaia and TESS frequencies. The black line indicates our chosen threshold for matching signals ($\Delta f<0.006$ d-1).
@@ 2.4 Comparison to Gaia frequencies
Figure 2 compares the main frequency from our TESS analysis with the frequencies derived from Gaia DR3 by Gaia Collaboration et al. (2023). As already found by Hey & Aerts (2024), most of the pulsation frequencies agree within a tolerance of $\Delta f<0.006$ d-1, highlighting once again the potential of the Gaia time series photometry to discover non-radial pulsators at a level of 4 mmag or above. For the twelve stars without a matching TESS primary frequency, we searched through other signals in order of decreasing amplitudes in the TESS data and compared them to the Gaia frequency. Once a match within our permitted tolerance was found we stopped the search. For seven stars [^1] The seventh star is not shown in Fig. 2 because its Gaia frequency $f_{\emph{Gaia}}=21$ d-1 is outside the $\beta$ Cep frequency range. Yet this star might still be a $\beta$ Cep pulsator as the dominant TESS frequency $f_{1}=6.6$ d-1., we found no matching signal in the TESS data. The median difference between the TESS and Gaia matched frequencies is $\Delta\tilde{f}=0.0013$ d-1.
We limited the current analysis to the primary frequency detected with Gaia. Although we possess measurements of the secondary frequency from the Gaia DR3 time series photometry, we noticed that these frequencies match our extracted frequencies only for a minority of cases. Often, we find $f_{\emph{Gaia}}>10$ d-1, whereas we rarely extracted any significant frequency in this range from TESS. It was already discussed by Gaia Collaboration et al. (2023) that the Gaia DR3 data suffer from instrumental effects at mmag level for the secondary frequencies, which may be connected to multiples of the satellite frequency of 4 d-1 for some of the variables. Hence for the $\beta$ Cep population study in our work, using the secondary frequency from Gaia is not informative. We note that detailed work on certain stars might be able to derive additional constraints for the pulsation mode of the secondary frequency. Further, longer time series data as delivered by future Gaia data releases might provide more robust secondary frequencies in agreement with TESS.
@@ 3.1 Gaia derived properties
We aim to analyse the $\beta$ Cep stars in our sample in a homogeneous way and hence chose to use stellar parameters from Gaia DR3. Of particular interest are the effective temperature ($T_{\mathrm{eff}}$) and the luminosity ($\log L/L_{\sun}$). Shi et al. (2023) show that the Gaia $T_{\mathrm{eff}}$ derived by the Extended Stellar Parametrizer for Hot Stars (ESP-HS, Fouesneau et al., 2023) is reliable for pulsating hot main sequence stars. We follow their example and use these values as the basis for our analysis.
In addition to the effective temperatures, we extracted positions, parallaxes, and $G$ band fluxes for every target. Where available, we also added stellar parameters from ESP-HS and the General Stellar Parametrizer from Photometry (GSPPHOT). Unfortunately, the Gaia ESP-HS parameters are not available for all stars in our sample but only for 190 of the 222 stars. Hence, we are not able to place all the observed stars in the Hertzsprung-Russell diagrams (HRD).
@@
The Gaia ESP-HS parameters also include the projected surface rotational velocity $v\sin i$ calculated from the line broadening of the RVS spectra under the assumption that rotation determines the line broadening, which is a valid assumption (cf., Aerts et al., 2023). Together with the luminosity and effective temperature, which gives an estimate of the stellar radius, we can calculate the surface rotation rate of the stars. This parameter, $f_{\mathrm{rot}}\sin i$, is available for 167 stars. Although, it is not very precise with a typical uncertainty $\sim 40\,\%$, it will still be useful in our subsequent analysis.
@@ 3.2 Gaia data of known $\beta$ Cep stars
To compare the properties of the Gaia discovered $\beta$ Cep stars with these of known $\beta$ Cep stars, we extracted their Gaia DR3 data as well. As the literature sample, we use the certain $\beta$ Cep stars from the catalogues of Stankov & Handler (2005), Pigulski & Pojmański (2008), Labadie-Bartz et al. (2020), Shi et al. (2024), and Eze & Handler (2024). For each of the catalogues, we derived the same Gaia properties as described above. From the four catalogues in the literature, we find 582 stars of which 408 are classified as genuine $\beta$ Cep stars. Among the 222 stars in the Gaia-discovered sample, 145 are not listed in any of the literature catalogues, bringing the total number of $\beta$ Cep stars in the combined sample to 553 stars.
@@ 4 Population of $\beta$ Cep stars
Before analysing our Gaia discovered sample in detail, we present the overall properties of the ensemble. Figure 3 shows all 442 stars (out of 553) with adequate Gaia ESP-HS parameters in an HRD. Most of the stars in our sample fall inside the instability strips for OB-type pulsators calculated by Burssens et al. (2020). These are based on one particular choice of input physics (relying on exponential core overshoot with parameter $f_{\mathrm{ov}}=0.02$) at solar metallicity so we expect some stars to occur outside that region, as also stressed by Aerts & Tkachenko (2024), Rehm et al. (2024), and Hey & Aerts (2024).
Figure 3: Hertzsprung-Russel diagram based on Gaia data of all 321 $\beta$ Cep stars from the literature (circles) and 121 new stars of our Gaia-informed sample (squares, new detections). The typical uncertainty is indicated in the lower left. As black outlines, we show the instability regions for p-modes (solid, $\beta$ Cep, $l=0-2$), and g-modes (dashed, SPB, $l=1-2$) based on the calculations from Burssens et al. (2020). In the background, we added in grey main sequence evolutionary tracks from Burssens et al. (2020). The colour-code gives the main pulsation frequency of the stars cut-off at 12 d-1. The top axis marks the spectral type.
Several stars from the literature occur in the colder slowly pulsating B star (SPB) instability region. We suspect that these stars were misclassified as $\beta$ Cep stars in the literature but are rather SPB stars. All the new Gaia-derived $\beta$ Cep stars fall inside or close to the $\beta$ Cep instability strip and can hence from their HRD position and TESS variability characterisation (see below) safely be assumed to be genuine $\beta$ Cep stars.
At $\log T_{\mathrm{eff}}\approx 4.5$, we find a gap between the bulk of the $\beta$ Cep stars and a smaller group mostly containing stars from our new sample. With only the Gaia ESP-HS data, we cannot be certain whether this gap is astrophysical or introduced by the intricate parameter estimation procedure. It is present in the entire ESP-HS data set. We note the position of the gap overlaps with the transition from spectral type O to B (Pecaut & Mamajek, 2013). In case the gap is physical these stars on the hotter side of the blue border of the strip would be of great interest for follow-up observations and detailed asteroseismic modelling as these stars might be binary or merger products (Fabry et al., 2022, 2023; Bellinger et al., 2024; Henneco et al., 2024).
@@ 4.1 Detailed analysis of the Gaia sample
After verifying that all of our sample stars with Gaia ESP-HS parameters fall indeed into the parameter range of $\beta$ Cep stars, we analyse their TESS light curves in more detail. In a visual inspection, three different light curve shapes emerge as the main sub-classes (Fig. 4). The first class comprises of mono-periodic light curves which show one dominant coherent oscillation over a TESS sector. On closer inspection after pre-whitening, these stars pulsate often in multiple frequencies as well but their secondary frequencies have small amplitudes compared to the dominant mode. Clearly multi-periodic pulsators constitute the second class. Their light curves exhibit modulations due to mode beating with periods between one and tens of days, as was already known from ground-based multisite campaigns prior to the space asteroseismology era. We distinguish between $\beta$ Cep multi-periodic light curves due to the presence of multiple low-order p- or g-modes and a third sub-class of hybrid pulsators with SPB components because their light curves have distinctly different shapes (Fig. 4).
@@
Figure 4: Examples of single sector TESS light curves of $\beta$ Cep variables of the different light curve sub-classes. The top light curve is a typical mono-periodic pulsator with a constant pulsation amplitude. The second and third rows show multi-periodic pulsators with different lengths of their beating patterns. For the third row the beating pattern is significantly longer than one TESS sector, hence, multiple TESS sectors are needed to resolve the close frequency peaks. The bottom row shows a typical SPB and $\beta$ Cep hybrid pulsator. Due to the very different time scales of the main pulsation modes its light curve has a distinct shape.
@@
We classified all 216 light curves based on their shape and the content of their periodogram into these three main classes. We find 49 (almost) mono-periodic, 126 multi-periodic, and 24 hybrid pulsators. The remaining 17 light curves show variability that does not fall into the three classes. Stankov & Handler (2005) classified the stars in their catalogue into similar categories and found a fraction 40 % of mono-periodic pulsators. However, they already expected this fraction to be lower due to unobserved low amplitude pulsations in the ground-based light curves, which have much higher noise level than the TESS space photometry. Our fraction of 24 % pulsators with one dominant oscillation mode is in agreement with this expectation.
Among the 17 stars we exclude from the $\beta$ Cep sub-classes, TIC 22370803, TIC 311328485, and TIC 460977279 are likely short period eclipsing binaries mimicking a pulsation signal in the right frequency domain. We remove these stars in the subsequent analysis. TIC 234644092 might also be of this class but without compelling evidence, we kept the star in our analysis. In contrast, TIC 285370131, TIC 450383484, TIC 329992262, TIC 399436828, and TIC 410100850 show $\beta$ Cep pulsations but their light curves are dominated by long-term quasi-periodic behaviour. We postulate that they are either Be stars or strongly contaminated. TIC 248275346 has additional high frequency components, which create a beating pattern unlike any other star in our sample. The three stars TIC 326815593, TIC 327282466, and TIC 365623418 are noisy and modulated with additional non-periodic variability which stems likely from flux contamination of neighbouring stars. TIC 217602397, TIC 464732844, and TIC 153657168 exhibit very noisy light curves with hidden periodic variability. Finally, TIC 406696133 is likely contaminated by a nearby, very red ($G-J=4.3$ mag) Gaia long-period variable. In the TESS observations this contaminant is brighter than our target star.
The stars TIC 90964721, TIC 395218466, and TIC 1861614014 are potentially (longer-period) eclipsing binaries with $\beta$ Cep pulsations, making them interesting targets for detailed modelling. The latter of the three is also included in the sample of Eze & Handler (2024).
@@
The 199 stars in the three sub-classes are found across the whole instability region. While most of the hybrid pulsators are close or inside the overlap region of the two instability strips, we find some notable exceptions. In particular the distinct group at high $T_{\mathrm{eff}}$ (see above), hosts two hybrid pulsators. A HRD with the stars colour-coded by the classifications is shown in the Appendix Fig. 18.

### v2

@@ 2.3 Extracted frequencies
Figure 1: Stacked representation of the extracted independent, significant TESS frequencies for 212 of the Gaia-detected $\beta$ Cep stars, sorted by the TESS frequency with the largest amplitude. The colour-scale indicates the amplitude of each detected signal on a logarithmic scale. This figure includes only the stars whose main frequency occurs in the plotted $\beta$ Cep range. Three stars have lower main frequencies and turn out to be SPB stars or $\beta$ Cep and SPB hybrid pulsators with dominant high-order g-modes, or are dominated by other variability than pulsations.
To analyse a large number of pulsators, stacking and sorting the periodograms can reveal hidden structures (e.g. Gilliland et al., 2010; Michel et al., 2017; Li et al., 2020b; Hey & Aerts, 2024; Read et al., 2024). Figure 1 shows the extracted significant frequencies from the TESS photometry for our Gaia-observed sample sorted by the frequency of highest amplitude. Since $\beta$ Cep stars show typically few excited low-order modes that do not occur in an asymptotic regime of high or low frequencies, no obvious structure emerges from this diagram. Nevertheless, this visual representation of the sample shows the range of detected significant frequencies. We observe that the distribution flattens towards low frequencies while declining more abruptly towards high frequencies. This is fully in line with instability computations of low-order modes in young and evolved $\beta$ Cep stars (Pamyatnykh, 1999). However, the sharp cut-off near 8 d-1 may also be partly due to the natural transition towards the $\delta$ Sct star frequency regime adopted in the classification algorithms (Gaia Collaboration et al., 2023). Further, it becomes obvious from the figure that the observed significant frequencies for a given star depend on the position of its dominant frequency.
Figure 2: Comparison between the highest amplitude Gaia and TESS frequencies for the 215 stars with significant TESS frequencies. Top: Dominant TESS frequency $f_{\mathrm{TESS}}$ against the Gaia derived frequency $f_{\emph{Gaia}}$. Black symbols indicate stars for which we found a match in frequency, blue symbols show stars matching a frequency other than the dominant signal, and red symbols indicate the seven stars for which the Gaia frequency does not match any frequency in the TESS data. Lines connect the offset frequencies to their matched counterpart. For three stars the main frequency is below the plotting range. Bottom: Difference between the Gaia and TESS frequencies. The black dashed line indicates our chosen threshold for matching signals ($\Delta f<0.006$ d-1).
@@ 2.4 Comparison to Gaia frequencies
Figure 2 compares the main frequency from our TESS analysis with the frequencies derived from Gaia DR3 by Gaia Collaboration et al. (2023). As already found by Hey & Aerts (2024), most of the pulsation frequencies agree within a tolerance of $\Delta f<0.006$ d-1, highlighting the potential of the Gaia time series photometry to discover non-radial pulsators at a level of 4 mmag or above. For the twelve stars without a matching TESS primary frequency, we searched through other signals in order of decreasing amplitudes in the TESS data and compared them to the Gaia frequency. Once a match within our permitted tolerance was found we stopped the search. For seven stars [^2] The seventh star is not shown in Fig. 2 because its Gaia frequency $f_{\emph{Gaia}}=21$ d-1 is outside the $\beta$ Cep frequency range. Yet this star might still be a $\beta$ Cep pulsator as the dominant TESS frequency $f_{1}=6.6$ d-1., we found no matching signal in the TESS data. The median difference between the TESS and Gaia matched frequencies is $\Delta\tilde{f}=0.0013$ d-1.
We limited the current analysis to the primary frequency detected with Gaia. Although we possess measurements of the secondary frequency from the Gaia DR3 time series photometry, we noticed that these frequencies match our extracted frequencies only for a minority of cases. Often, we find $f_{\emph{Gaia}}>10$ d-1, whereas we rarely extracted any significant frequency in this range from TESS. It was already discussed by Gaia Collaboration et al. (2023) that the Gaia DR3 data suffer from instrumental effects at mmag level for the secondary frequencies, which may be connected to multiples of the satellite frequency of 4 d-1. Hence for the $\beta$ Cep population study in our work, using the secondary frequency from Gaia is not informative. We note that detailed work on certain stars might be able to derive additional constraints for the pulsation mode of the secondary frequency. Further, longer time series data delivered by future Gaia data releases might provide more robust secondary frequencies in agreement with TESS.
@@ 3.1 Gaia derived properties
We aim to analyse the $\beta$ Cep stars in our sample in a homogeneous way and hence chose to use stellar parameters from Gaia DR3. Of particular interest are the effective temperature ($T_{\mathrm{eff}}$) and the luminosity ($\log L/L_{\sun}$). Shi et al. (2023) show that the Gaia $T_{\mathrm{eff}}$ derived by the Extended Stellar Parametrizer for Hot Stars (ESP-HS, Fouesneau et al., 2023) is reliable for pulsating hot main sequence stars. We follow their example and use these values as the basis for our analysis. Further, we extracted positions, parallaxes, and $G$ band magnitudes for every target. Where available, we also added stellar parameters from ESP-HS and the General Stellar Parametrizer from Photometry (GSPPHOT). Unfortunately, the Gaia ESP-HS parameters are only available for 190 of the 222 stars. Hence, we are not able to place all the observed stars in the Hertzsprung-Russell diagrams (HRD). Further, this limits the sample of stars that we can model asteroseismically.
@@
An estimate of the projected surface rotation frequency is useful when searching for rotationally split multiplets and examining differential rotation. The Gaia ESP-HS parameters also include the projected surface rotational velocity $v\sin i$ calculated from the line broadening of the RVS spectra under the assumption that rotation determines the line broadening, which is a valid assumption (cf., Aerts et al., 2023). Together with the luminosity and effective temperature, which gives an estimate of the stellar radius, we can calculate the projected surface rotation rate of the stars. This parameter, $f_{\mathrm{rot}}\sin i$, is available for 167 stars. Although it is not very precise with a typical uncertainty $\sim 40\,\%$, it still proves useful in our subsequent analysis.
@@ 3.2 Gaia data of known $\beta$ Cep stars
To compare the properties of the Gaia-detected $\beta$ Cep stars with these of known $\beta$ Cep stars, we extracted their Gaia DR3 data as well. As the literature sample, we use the genuine $\beta$ Cep stars from the catalogues of Stankov & Handler (2005), Pigulski & Pojmański (2008), Labadie-Bartz et al. (2020), Shi et al. (2024), and Eze & Handler (2024). For each of the catalogues, we derived the same Gaia properties as described above. From the four catalogues in the literature, we find 582 stars of which 408 are classified as genuine $\beta$ Cep stars. Among with the 222 $\beta$ Cep candidates in our sample – 145 are not listed in any of the literature catalogues – the total number of $\beta$ Cep stars in the combined sample is 553 stars.
@@ 4 Population of $\beta$ Cep stars
Before analysing our Gaia-detected sample in detail, we present the overall properties of the ensemble. Figure 3 shows all 442 stars (out of 553) with adequate Gaia ESP-HS parameters in an HRD. Most of the stars in our sample fall inside the instability strips for OB-type pulsators calculated by Burssens et al. (2020). These are based on one particular choice of input physics (relying on exponential core overshoot with parameter $f_{\mathrm{ov}}=0.02$) at solar metallicity, so we expect some stars to occur outside that region, as also stressed by Aerts & Tkachenko (2024), Rehm et al. (2024), and Hey & Aerts (2024).
Figure 3: Hertzsprung-Russell diagram based on Gaia data of all 321 $\beta$ Cep stars from the literature (circles) and 121 new stars of our Gaia-informed sample (squares, new detections). The typical uncertainty is indicated in the lower left. Black outlines show the instability regions for p-modes (solid, $\beta$ Cep, $l=0-2$), and g-modes (dashed, SPB, $l=1-2$) based on the calculations from Burssens et al. (2020). Main sequence evolutionary tracks of Burssens et al. (2020) are shown in grey. The colour-code gives the main pulsation frequency of the stars cut-off at 12 d-1. Stars marked with a red cross do not exhibit $\beta$ Cep variability in TESS. The top axis marks the spectral type.
Several stars from the literature occur below the $\beta$ Cep instability region. We checked these stars with publicly available light curves from the TESS Quick Look Pipeline (QLP, Huang et al., 2020) using a classification similar to our main sample and outlined below in Sect. 4.1. We find the majority of stars close to the instability region to be $\beta$ Cep stars pointing either to imprecisions of the stellar parameters or a larger extend of the instability region as found for other pulsators in this area of the HRD too (Gaia Collaboration et al., 2023; Mombarg et al., 2024). Stars further away from the instability region are either short-period eclipsing binary stars or SPB stars. The stars [^3] TIC 17449899 (V2187 Cyg), TIC 23548300 (HD 55708), TIC 129364361 (NGC 2483 2), TIC 130826839 (HD 108769), TIC 155571411 (HD 69016), TIC 312931751 (HD 116827), TIC 337165095 (HD 42401), and TIC 427406031 (HD 143605) with obvious non-$\beta$ Cep variability (and available light curves) are marked in Fig. 3 with red crosses. All the new Gaia-derived $\beta$ Cep stars fall inside or close to the $\beta$ Cep instability strip and can hence safely be assumed to be genuine $\beta$ Cep stars from their HRD position and TESS variability characterisation (see below).
At $\log T_{\mathrm{eff}}\approx 4.5$, we find a gap between the bulk of the $\beta$ Cep stars and a smaller group mostly containing stars from our new sample. With only the Gaia ESP-HS data, we cannot be certain whether this gap is astrophysical or introduced by the intricate parameter estimation procedure. It is present in the entire ESP-HS data set. We note the position of the gap overlaps with the transition from spectral type O to B (Pecaut & Mamajek, 2013). In case the gap is physical the stars on its hotter side would be of great interest for follow-up observations and detailed asteroseismic modelling as these stars might be products of binary evolution (Fabry et al., 2022, 2023).
@@ 4.1 Detailed analysis of the Gaia sample
After verifying that all of our sample stars with Gaia ESP-HS parameters fall indeed into the parameter range of $\beta$ Cep stars, we analyse their TESS light curves in more detail. In a visual inspection of the light curves and their Fourier transforms, three main sub-classes emerge (Fig. 4). The first class comprises of mono-periodic pulsators which show one dominant coherent oscillation in the entire light curve. The periodograms of these stars often reveal they pulsate in multiple frequencies, but their secondary frequencies have small amplitudes compared to the dominant mode. Clearly multi-periodic pulsators constitute the second class. Their light curves exhibit modulations due to mode beating with periods between one and tens of days, as already known from ground-based multisite campaigns prior to the space asteroseismology era. In some cases, the beating pattern is significantly longer than one TESS sector. Hence, multiple TESS sectors are needed to resolve the close frequency peaks. We distinguish between $\beta$ Cep multi-periodic light curves (due to the presence of multiple low-order p- or g-modes) and hybrid pulsators with SPB components as a third sub-class because their secondary signals have a much smaller frequency, resulting in a differently shaped light curve. (Fig. 4). We manually classified all 216 light curves based on their shape and the content of their periodogram into these three main classes.
@@
Figure 4: Examples of single sector TESS light curves (left) of $\beta$ Cep variables of the different light curve sub-classes and their associated periodograms (right). The top light curve is a typical mono-periodic pulsator with a constant pulsation amplitude and a single peak in the frequency spectrum. The second and third panels show multi-periodic pulsators with different lengths of their beating patterns. The bottom panel shows a typical $\beta$ Cep and SPB hybrid pulsator.
In addition to the manual classification, we quantified the light curve shapes by calculating two measures of the amplitude variability. Firstly, we use the length of the main beating pattern given by
$P_{\mathrm{beat}}=\left|f_{1}-f_{2}\right|^{-1}.$ (1)
Secondly, we calculate the strength of amplitude variability
$A_{\mathrm{var}}=A_{90}/A_{95}\cdot 0.95/0.90,$ (2)
in which $A_{95}$ ($A_{90}$) is the difference between the 95th (90th) and 5th (10th) percentile of the light curve. The percentile differences ($A_{95}$) are known to robustly estimate stellar variability in space photometry (Basri et al., 2011). We expand on this idea and use the ratios of the percentile differences. By normalising this ratio, a pure sine variability (i.e. a mono-periodic pulsator) reaches unity, while multiperiodic stars can be found at lower values. We used the full light curve itself to compute $A_{\mathrm{var}}$ rather than the amplitudes of the extracted frequencies because it includes all variability beyond the two main frequencies in this measure.
Figure 5 shows that we can clearly separate the above defined groups based on $A_{\mathrm{var}}$ and the beating period. The initial manual classifications are mostly confirmed. Mono-periodic stars fall towards the right. Multi-periodic pulsators populate the upper-left because they are not close to a single sinusoid, yet have a long beating pattern, while hybrids lie near the lower-left because their primary and secondary frequencies differ greatly. Some stars fall into regimes in Fig. 5 that are not associated with their manual classification. Among the mono-periodic stars, several stars display additional rotational variability that modulated the light curve without creating a beating pattern, a feature that cannot be accounted for with the current parameters. Some stars have long beating patterns with low amplitudes which makes them appear mono-periodic. The grouping of mono-periodic stars at long beating periods indicate that these stars are indeed not truly mono-periodic but exhibit multi-periodicity with very small frequency differences, a well-known phenomenon of some classical $\beta$ Cep stars (e.g. HD 129929, $\beta$ CMa). Given the few differences between our manual classification and the regimes in Fig. 5, we provide $A_{\mathrm{var}}$ and $P_{\mathrm{beat}}$ in Table 1 to enable quantified distinctions of the different groups.
Figure 5: Beating period of the main two frequencies ($P_{\mathrm{beat}}$) against the normalised percentile amplitude ratio ($A_{\mathrm{var}}$). Light curves manually classified as mono-periodic are shown in purple, multi-periodic in olive, hybrids in red and other cases in cyan. The dashed lines indicate the approximate dividing lines of the groups of pulsating stars. Uncertainties on the beating period are within the symbol sizes ($A_{\mathrm{var}}$ has no formal uncertainty due to its definition).
@@
Based on the manual classification, we find 49 (almost) mono-periodic, 126 multi-periodic, and 24 hybrid pulsators. The remaining 17 light curves show variability that does not fall into the three classes. Using $A_{\mathrm{var}}$ and $P_{\mathrm{beat}}$ the above introduced quantification for the number of multi-periodic stars stays constant at 124, while we find fewer true mono-periodic stars (33) and more hybrid pulsators (42). Stankov & Handler (2005) classified the stars in their catalogue into similar categories and found a fraction 40 % of mono-periodic pulsators. However, they expected this fraction to be lower due to unobserved low amplitude pulsations in the ground-based light curves, which have much higher noise level than the TESS space photometry. Our fraction of 24 % (16.5 % when using the classification in Fig. 5) of pulsators with one very dominant oscillation mode is in agreement with this expectation.
@@
The stars of the three sub-classes are found across the whole instability region in the HRD. While most of the hybrid pulsators are close to or inside of the overlap region of the two instability strips, we find some notable exceptions. In particular the distinct group at high $T_{\mathrm{eff}}$ (see above), hosts two hybrid pulsators. A HRD with the stars colour-coded by the classifications is shown in the Appendix Fig. 21.
Among the 17 stars we exclude from the $\beta$ Cep sub-classes, three stars (TIC 22370803, TIC 311328485, and TIC 460977279 (HD 137405)) are likely short period eclipsing binaries mimicking a pulsation signal in the right frequency domain. We remove these stars in the subsequent analysis. TIC 234644092 might also be of this class but without compelling evidence, we kept the star in our analysis. In contrast, TIC 285370131, TIC 450383484, TIC 329992262 (V1155 Cas), TIC 399436828, and TIC 410100850 show $\beta$ Cep pulsations but their light curves are dominated by long-term quasi-periodic behaviour. We postulate that they are either Be stars or strongly contaminated. TIC 248275346 (HD 159792) has additional high frequency components, which create a beating pattern unlike any other star in our sample. The three stars TIC 326815593, TIC 327282466, and TIC 365623418 are noisy and modulated with additional non-periodic variability which stems likely from flux contamination of neighbouring stars. TIC 217602397, TIC 464732844, and TIC 153657168 (HD 114733) exhibit very noisy light curves with hidden periodic variability. Finally, TIC 406696133 is likely contaminated by a nearby, very red ($G-J=4.3$ mag) Gaia long-period variable. In the TESS observations this contaminant is brighter than our target star.
The stars TIC 90964721, TIC 395218466, and TIC 1861614014 (HD 339003) are potentially (longer-period) eclipsing binaries with $\beta$ Cep pulsations, making them interesting targets for detailed modelling. The latter of the three is also included in the sample of Eze & Handler (2024).